) Suppose that your business is operating at the four-and-a-half-sigma quality level (with 1350 ppm). If projects have an average improvement rate of 6%monthly, how many years will it take to achieve six-sigma quality (with 3.4 ppm)?
Answer #1

AT Six Sigma, there will be 3.4 ppm

Currently there is 1350 ppm

Let assume in x month, business will achieve Six Sigma quality.

3.4 = 1350*(1-6%)^x

0.94^x = 3.4/1350

x = ln(3.4/1350)/ln(0.94)

x = 96.71 month

x = 8.06 years